please show steps to solve: (1-x)² +(1/3)²=x²
Answer by CharlesG2(834)

please show steps to solve: (1-x)² +(1/3)²=x²
(1-x)² +(1/3)²=x²
(1-x)(1-x) + (1/3)(1/3) = x^2
1 - 2x + x^2 + 1/9 = x^2
10/9 - 2x = 0 (the x^2 cancel out)
-2x = -10/9
2x = 10/9
x = 5/9